OpenStudy (anonymous):

The sum of two numbers is 20. Find the numbers and their product if their product is a maximum.

OpenStudy (badhi):

x+y=20 C=xy =x(20-x) =20x-x^2 if you know calculus, dC/dx=20-2x=0 x=10 (when maximum) y=10 if not, C=-(x-10)^2+100 C becomes maximum when x=10 (-(x-10)^2>0) x=10 when maximum y=10

OpenStudy (anonymous):

thanks i get it... i'll do it the noncalculus way XD